well this is actually from algebra 2.
3 The characteristic of log 6198 is _______
2
4
3
1

i really really have no idea how to do this problem... please help!

Answer by venugopalramana(3286)

well this is actually from algebra 2.
3 The characteristic of log 6198.....WRITE THE GIVEN NUMBER AS X*10^N
WHERE X IS A NUMBER WITH ONLY UNITS PLACE AND THEN N IS THE CHARACTERSTIC.HERE
6198 = 6.198*1000= 6.198*10^3...HENCE 3 IS THE CHARACTERSTIC


is _______
2
4
3
1